OpenStudy (anonymous):

12x[1]+3x[2]=9 -8x[1]-2x[2]=-6

hero (hero):

Use ^ if you're talking about exponents. For example, x squared = x^2

OpenStudy (anonymous):

yes but they are at the bottom of the equation so like 12x\[_{1}\]

hero (hero):

Okay, so use x and y instead

hero (hero):

If x[1] = x and x[2] = y, then: 12x + 3y = 9 -8x - 2y = -6 Solve the system

hero (hero):

You should know how to solve it now.